OpenStudy (anonymous):

3a ti the second power -2 a=5

OpenStudy (anonymous):

times by the second power

OpenStudy (anonymous):

\[3a^2-2a=5?\]

OpenStudy (anonymous):

no 5 is what (a) is

OpenStudy (anonymous):

so 3a^2-2 and a=5

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok, since a=5 just put that in where a is. 3(5)^2-2 and solve

OpenStudy (anonymous):

so would the answer be 43

OpenStudy (anonymous):

Remember PEMDAS? First square the 5. 3*25-2 then multiply 75-2 then subtract 73

OpenStudy (anonymous):

i thought you would do the exponents 3a*3a=9a plug in a so 9*5=45-2=43

OpenStudy (anonymous):

No, the equation is 3a^2-2 You're given a=5 so just put that into the equation where a is. 3(5)^2-2 and that gives you 73

OpenStudy (anonymous):

i asked if id have to do 15*15 and he said no